OpenStudy (anonymous):

solve each equation... x^4-x=0 3x^4+18=21x^2 2x^4-26x^2-28=0 5x^4+50x^2+80=0 helpp pleasee!!!!!!!

OpenStudy (campbell_st):

for the 1st one take x as a common factor so you have x(x^3 -1) = 0 so x = 0 or 1

OpenStudy (kira_yamato):

3x^4+18=21x^2 3x^4-21x^2+18=0 Let y = x^2 3y^2 - 21y + 18 = 0 y = -1 or -6 x = ±i or ±(sqrt(6))i

OpenStudy (kira_yamato):

2x^4-26x^2-28=0 x^4-13x^2-14=0 Let y = x^2 y^2 - 13y -14=0 y = -1 or 14 x = ±i or ±sqrt(14)

OpenStudy (kira_yamato):

5x^4 + 50x^2 + 80 = 0 x^4 + 10x^2 + 16 = 0 Let y = x^2 y^2 + 10y + 16 = 0 y = -2 or -8 x = (sqrt(2))i or (2sqrt(2))i